# Differentially Private Random Feature Model

### v1

@@ Abstract
Designing privacy-preserving machine learning algorithms has received great attention in recent years, especially in the setting when the data contains sensitive information. Differential privacy (DP) is a widely used mechanism for data analysis with privacy guarantees. In this paper, we produce a differentially private random feature model. Random features, which were proposed to approximate large-scale kernel machines, have been used to study privacy-preserving kernel machines as well. We consider the over-parametrized regime (more features than samples) where the non-private random feature model is learned via solving the min-norm interpolation problem, and then we apply output perturbation techniques to produce a private model. We show that our method preserves privacy and derive a generalization error bound for the method. To the best of our knowledge, we are the first to consider privacy-preserving random feature models in the over-parametrized regime and provide theoretical guarantees. We empirically compare our method with other privacy-preserving learning methods in the literature as well. Our results show that our approach is superior to the other methods in terms of generalization performance on synthetic data and benchmark data sets. Additionally, it was recently observed that DP mechanisms may exhibit and exacerbate disparate impact, which means that the outcomes of DP learning algorithms vary significantly among different groups. We show that both theoretically and empirically, random features have the potential to reduce disparate impact, and hence achieve better fairness.
@@ 1 Introduction
Over the last decades, data privacy has become a central concern due to the great amount of personal data, such as medical records, stored and used in ubiquitous applications. Machine learning algorithms have been shown to have the ability to explore population level patterns, however the outputs of machine learning algorithms may reveal sensitive personal information. Thus, one of the challenges for machine learning algorithms is how to protect the privacy of sensitive personal information in the training data. Towards this end, ($\epsilon,\delta$)-differential privacy (DP) [1] has emerged as a gold standard, as it offers formal and quantitative privacy guarantees and enjoys many attractive properties from an algorithmic design perspective. For example, differential privacy protects against arbitrary risks, beyond just protection against re-identification. Moreover, differential privacy satisfies composition rules. For ”sequential composition”, if two mechanisms $\mathcal{M}$ and $\mathcal{M}^{\prime}$ satisfy $(\epsilon_{1},\delta_{1})$-DP and $(\epsilon_{2},\delta_{p})$-DP, respectively, then the sequential model $\mathcal{M}^{*}=(\mathcal{M},\mathcal{M}^{\prime})$ is $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2})$-DP. For ”parallel composition”, if mechanisms $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP and $\{X_{1},\dots,X_{k}\}$ are distinct datasets, then the parallel mechanism $\mathcal{M}^{*}(X)=(\mathcal{M}(X_{1}),\dots,\mathcal{M}(X_{k}))$ also satisfies $(\epsilon,\delta)$-DP. The composition of a data-independent mapping of $f$ with an $(\epsilon,\delta)$-DP is also $(\epsilon,\delta)$-DP, which is also called immune to post-processing. Informally speaking, a randomized mechanism is differentially private if its (distribution of) outcomes remain roughly the same by changing or removing any single individual in a dataset.
@@
In this paper, we focus on the random feature model, which was proposed to approximate large-scale kernel machines [9, 10]. In addition, random feature models refer to two-layer networks where the single hidden layer is randomized instead of being trained. Since only the weights in the output layer are trainable, it can be viewed as a linear approximation of a neural network. The training of random feature models reduces to solving a convex optimization problem. Random feature models have wide applications in regression problems [11], classification problems [12], dynamical systems [13], and scientific computing [14, 15, 16]. Therefore, exploring applications and examining the theoretical properties of random feature models provides a more profound comprehension of kernel machines and valuable insights into extremely wide or deep neural networks.
The random feature model is widely used to study differentially private learning with kernel methods, see [17, 18, 19, 20]. Linear models, such as support vector machines (SVM), can solve nonlinear cases by utilizing a kernel function. The idea of kernel method is mapping features to a Hilbert space where the inner product can be computed by a kernel function. However, kernel-based models have a fundamental weakness in achieving differential privacy. According to the celebrated representer theorem [21], the optimal model is given by a linear combination of kernel functions centered at the data points. Thus, even if the model is computed in a privacy-preserving way, it also reveals the training data. The random feature method, which approximates the kernel function using random projections, is used to avoid the aforementioned problem.
@@ 1.1 Contributions
The goal of this paper is to construct differentially private random feature models along with privacy and generalization guarantees. We put ourselves in the over-parametrized regime where the number of random features is greater than the number of samples. We consider the random feature models obtained by solving the min-norm interpolation problem. To the best of our knowledge, it is the first work on the differentially private min-norm interpolation random feature model. This setting is of great interest given that neural networks always have a large number of parameters and the interpolation estimators can generalize well, see [22, 23, 24]. Following the output perturbation mechanism, we propose Algorithm 1 that produces a private random feature model. This algorithm comes with theoretical guarantees for differential privacy and generalization error. Moreover, we show that the proposed DP random feature model has the potential to reduce disparate impact. More precisely, our contributions are as follows:
@@
    We propose a differentially private random feature model (Algorithm 1) based on min-norm interpolator and derive its theoretical guarantees for differential privacy and generalization error with high probability, see Section 3. Our result holds without the need of regularization and relies on the concentration property of the random feature matrix.
@@
    Our experimental results in Section 4 on synthetic data and well-known datasets demonstrate that our proposed algorithm achieves competitive generalization performance when compared to the state-of-the-art DP random feature methods in [20, 25].
@@
    We also show that the proposed DP random feature model does not suffer disparate impact and achieve better fairness both theoretically in Section 3 and empirically in Section 4.
@@ 1.2 Related work
DP random feature model. There is a large body of literature on privacy preserving kernel-based machine learning algorithms [17, 18, 19, 20, 25]. In [19, 18], the authors studied a private kernel Support Vector Machine (SVM) via random feature approximation of a shift-invariant kernel. In [17, 20], the authors studied the problem of privacy preserving learning using empirical risk minimization and then applied their results to produce privacy-preserving kernel-based models. All previous work considered the regularization regime where a strongly convex regularization term ($\ell_{2}$ regularization for example) is added to the empirical risk function. Their theoretical guarantees depend on the choice of regularization parameter, which cannot be generalized to the interpolation regime. In [25], the authors focused on the differentially private stochastic gradient descent (SGD) algorithm based on random features. The theoretical guarantees rely on technical assumptions on the kernel and a certain spectral decay rate on the second moment operator, which may be difficult to verify in practice. The obtained convergence rate depends on the sample size, which may be hard to obtain since increasing the sample size is hard in real-world applications.
Fairness. Designing private fair learning algorithms has been studied in prior work [3, 2, 6, 5, 26, 27, 28]. Focusing on the classification setting, it was shown that the performance impact of privacy may be more severe on underrepresented groups, resulting in accuracy disparity [2]. In [3], the authors studied how group sample size and group clipping bias affect the impact of differential privacy in DP stochastic gradient descent and how adaptive clipping for each group helps to mitigate the disparate impact. The trade-offs between privacy and fairness were studied in [6, 27] in the regression setting. Stratification has proven to be a simple but powerful technique to remove disparate impact. It has been studied for mean estimation in [28] and binary classification in [29].
The remainder of the paper is organized as follows. In Section 2, we present preliminaries on the random feature model and differential privacy. Our main algorithm and its theoretical guarantees are introduced in Section 3. In Section 4, we empirically explore the trade-offs achieved by our post-processing algorithm on synthetic datasets and real datasets. We conclude the paper with several future directions in Section 5.
@@ 2.1 Notation
In this paper, we let $\mathbb{R}$ be the set of all real numbers and $\mathbb{C}$ be the set of all complex numbers where $i=\sqrt{-1}$ denotes the imaginary unit. We define the set $[N]$ to be all natural numbers smaller than or equal to $N$, i.e. $\{1,2,\dots,N\}$. We use bold letters to denote vectors or matrices, and denote the identity matrix of size $n\times n$ by $\mathbf{I}_{n}$. For any two vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{d}$, the inner product is denoted by $\langle\mathbf{x},\mathbf{y}\rangle=\sum_{j=1}^{d}x_{j}\bar{y}_{j}$, where $\mathbf{x}=[x_{1},\dots,x_{d}]^{\top}$ and $\mathbf{y}=[y_{1},\dots,y_{d}]^{\top}$. For a vector $\mathbf{x}\in\mathbb{C}^{d}$, we denote by $\|\mathbf{x}\|_{p}$ the $\ell_{p}$-norm of $\mathbf{x}$. For a matrix $\mathbf{A}\in\mathbb{C}^{m\times N}$, its conjugate transpose is denoted by $\mathbf{A}^{*}$. If $\mathbf{A}\in\mathbb{C}^{m\times N}$ is full rank, then its Moore-Penrose inverse is $\mathbf{A}^{\dagger}=\mathbf{A}(\mathbf{A}\mathbf{A}^{*})^{-1}$ if $m\leq N$ or $\mathbf{A}^{\dagger}=(\mathbf{A}^{*}\mathbf{A})^{-1}\mathbf{A}^{*}$ if $m\geq N$. The (induced) $p$-norm of matrix $\mathbf{A}$ is written as $\|\mathbf{A}\|_{p}$.
2.2 Random Feature Model
Theorem 1 (Bochner [30]).
[…]
The kernel function can be approximated by $k(\mathbf{x},\mathbf{x}^{\prime})\approx\langle\phi(\mathbf{x}),\phi(\mathbf{x}^{\prime})\rangle$. To produce feature map $\phi$ mapping data to finite-dimensional space $\mathbb{R}^{N}$, we can use random cosine features [31]. Setting random feature $\boldsymbol{\omega}\in\mathbb{R}^{d}$ generated from $\rho(\boldsymbol{\omega})$ and $b\in\mathbb{R}$ sampled from uniform distribution on $[-\pi,\pi]$, the random cosine feature is defined as $\phi(\mathbf{x},\boldsymbol{\omega},b)=\cos(\langle\boldsymbol{\omega},\mathbf{x}\rangle+b)$. For the sake of simplicity, we consider the random Fourier features in this paper. The results can be generalized to random cosine features, and therefore approximate real-valued function $f:\mathbb{R}^{d}\to\mathbb{R}$.
then problem (3) is also called the ridgeless regression problem.
@@ 3.2 Privacy Guarantee
Theorem 5 (Concentration of Random Feature Matrix).
@@
Remark.
Notice that our privacy result depends on the concentration properties of the random feature matrix $\mathbf{A}$, which has been studied in [32]. The conditions on the concentration result can be adopted to some other settings, for example, when the data samples $\{\mathbf{x}_{j}\}_{j\in[m]}$ are well separated and the random features $\{\boldsymbol{\omega}_{k}\}_{k\in[N]}$ are sampled from sub-gaussian distribution.
@@ 3.3 Generalization Error
In this section, we will derive an upper bound for the generalization error of our proposed differentially private random feature model. Let $f:\mathbb{R}^{d}\to\mathbb{C}$ be the target function belonging to the function class
@@
where $\rho(\boldsymbol{\omega})$ is a probability distribution defined on $\mathbb{R}^{d}$. Function space is indeed a Reproducing kernel Hilbert space with associated kernel function defined in (2) via random Fourier features.
@@
[…]
where $\mu$ is defined by the sampling measure for the data $\mathbf{x}$. The accuracy guarantee is provided below:
Theorem 7 (Generalization Error Bound for Private Random Feature Model).
@@
and bound each term separately. The first term is the standard generalization error of random feature models which appears in recent work [33, 24, 34, 35]. The second term characterizes the difference between non-private and private random feature models, which can be bounded by applying the sub-exponential tail bound. ∎
The generalization error bound of (non-private) random feature models has of great interests, see recent work [33, 24, 34, 35]. Our analysis follows the structure in [24, 33], where the authors provided an upper bound for generalization error using the concentration property of random feature matrix. However, their results only hold for functions $f$ belonging to a subset of $\mathcal{F}(\rho)$. In our analysis, we take a different ”best random feature approximation” $f^{*}$, which allows us to approximate all functions in $\mathcal{F}(\rho)$. Specifically, we define the best random feature approximation $f^{*}(\mathbf{x})$ as
@@
$\alpha_{\leq T}(\boldsymbol{\omega})=\begin{cases}\alpha(\boldsymbol{\omega})\quad&\mbox{ if }\left|\alpha(\boldsymbol{\omega})\right|\leq T\\ 0\quad&\mbox{ otherwise }\end{cases}\quad\mbox{ and }\quad\alpha_{>T}(\boldsymbol{\omega})=\begin{cases}\alpha(\boldsymbol{\omega})\quad&\mbox{ if }\left|\alpha(\boldsymbol{\omega})\right|>T\\ 0\quad&\mbox{ otherwise }.\end{cases}$
@@
Theorem 8 (Generalization Error Bound for Non-private Random Feature Model.).
@@
The first term is the approximation error which measures the difference between the target function and the best random feature model. The second term is the estimation error. To bound the second term, we utilize the concentration property of Random Feature matrix (Theorem 5) as well. ∎
We now bound the difference between the private random feature model and the non-private random feature model. Notice that the difference between the non-private coefficient vector $\mathbf{c}^{\sharp}$ and the private coefficient vector $\hat{\mathbf{c}}$ is indeed $\mathbf{z}\in\mathbb{R}^{N}$, which is generated from multivariate normal distribution $\mathcal{N}(0,\sigma^{2}\mathbf{I}_{N})$ with the variance $\sigma^{2}=\frac{8\ln(1.25/\delta_{p})}{N\epsilon_{p}^{2}(1-2\eta)}$ for some $\eta\in(0,1)$.
@@
$v(\mathbf{u}_{1},\dots,\mathbf{u}_{m})$ $:=\|f^{\sharp}-\hat{f}\|^{2}_{L^{2}(\mu)}-\frac{1}{m}\sum_{j=1}^{m}\left|f^{\sharp}(\mathbf{u}_{j})-\hat{f}(\mathbf{u}_{j})\right|^{2}$
@@
$\leq$ $\frac{2N}{m}\|\mathbf{c}^{\sharp}-\hat{\mathbf{c}}\|_{2}^{2}\leq\frac{2N}{m}\|\mathbf{z}\|_{2}^{2}:=\Delta_{v},$
@@
where the last inequality holds with probability at least $1-3\delta$ using Theorem 5. we apply Lemma 10 to estimate $\|\mathbf{z}\|_{2}^{2}$. Taking the square root for both sides leads to the desired bound.
@@
Notice that $\mathbf{z}_{k}^{2}$ is a sub-exponential random variable with parameters $(\nu_{k},\alpha_{k})=(2\sigma^{2},4\sigma^{2})$, then $\|\mathbf{z}\|_{2}^{2}$ is also a sub-exponential random variable with parameters $\nu=\sqrt{\sum_{k=1}^{N}\nu_{k}^{2}}$ and $\alpha=\alpha_{k}$. Then, we apply standard sub-exponential tail bound, see [36, Proposition 2.9], to obtain the desired estimation of $\|\mathbf{z}\|_{2}^{2}$. ∎
The output perturbation mechanisms were also employed in [20, 37], where the added random noise $\mathbf{z}$ follows a probability distribution with density function $\propto\exp(-\xi\|b\|_{2})$. To sample random variable $\mathbf{z}$ with density function $\propto\exp(-\xi\|b\|_{2})$, it suffices to i) sample the radius $R$ from Gamma($N$,$\eta$), then ii) sample $Y$ uniformly at random from the $\ell_{2}$-sphere of radius 1, and set $\mathbf{z}=RY$. Moreover, we can show that $\|\mathbf{z}\|_{2}$ follows a Gamma distribution Gamma($N$,$\xi$) where $\xi$ is selected as $\xi=\epsilon_{p}/\Delta(\mathbf{c}^{\sharp})$ to guarantee $(\epsilon_{p},0)$-differential privacy. By Chebyshev’s inequality, we have that

### v2

@@ Abstract
Designing privacy-preserving machine learning algorithms has received great attention in recent years, especially in the setting when the data contains sensitive information. Differential privacy (DP) is a widely used mechanism for data analysis with privacy guarantees. In this paper, we produce a differentially private random feature model. Random features, which were proposed to approximate large-scale kernel machines, have been used to study privacy-preserving kernel machines as well. We consider the over-parametrized regime (more features than samples) where the non-private random feature model is learned via solving the min-norm interpolation problem, and then we apply output perturbation techniques to produce a private model. We show that our method preserves privacy and derive a generalization error bound for the method. To the best of our knowledge, we are the first to consider privacy-preserving random feature models in the over-parametrized regime and provide theoretical guarantees. We empirically compare our method with other privacy-preserving learning methods in the literature as well. Our results show that our approach is superior to the other methods in terms of generalization performance on synthetic data and benchmark data sets. Additionally, it was recently observed that DP mechanisms may exhibit and exacerbate disparate impact, which means that the outcomes of DP learning algorithms vary significantly among different groups. We adopt excessive risk gap and statistical parity as the fairness metrics. Our results suggest, both theoretically and numerically, that random features have the potential to achieve better fairness.
@@ 1 Introduction
Over the last decades, data privacy has become a central concern due to the great amount of personal data, such as medical records, stored and used in ubiquitous applications. Machine learning algorithms have been shown to have the ability to explore population level patterns; however, the outputs of machine learning algorithms may reveal sensitive personal information. Thus, one of the challenges for machine learning algorithms is how to protect the privacy of sensitive personal information in the training data. Towards this end, ($\epsilon,\delta$)-differential privacy (DP) [1] has emerged as a gold standard. Essentially, a randomized mechanism is differentially private if its (distribution of) outcomes remain roughly the same by changing or removing any single individual in a dataset.
@@
In this paper, we study the random features model under the differentially private setting. The random features model, which was proposed to approximate large-scale kernel machines [9, 10], has wide applications in regression problems [11, 12], classification problems [13], signal processing [14], and scientific computing [15, 16, 17, 18]. In the context of differential privacy, the random feature model is used to study differentially private learning with kernel methods, see [19, 20, 21, 22]. Linear models, such as support vector machines (SVM), can solve nonlinear cases by utilizing a kernel function. The idea of kernel method is to map features to a Hilbert space where the inner product can be computed by a kernel function. However, kernel-based models have a fundamental weakness in achieving differential privacy. According to the celebrated representer theorem [23], the optimal model is given by a linear combination of kernel functions centered at the data points. Thus, even if the model is computed in a privacy-preserving way, it also reveals the training data. The random features method, which approximates the kernel function using random projections, can avoid the aforementioned problem.
In addition to approximating kernel machines, the random features model can be viewed as a two-layer network where the single hidden layer is randomized then fixed and only the weights in the output layer are trainable. The training of the random features model reduces to solving a convex optimization problem. Recently, a several papers [24, 25, 26, 27] have studied the training dynamics of over-parametrized neural networks (where the network size is larger than the sample size) through neural tangent kernel (NTK). The NTK theory showed that the trainable parameter of over-parametrized neural networks is close to its initialization, and hence the training dynamics of over-parametrized neural networks can be studied by over-parametrized random features. Therefore, exploring applications and examining the theoretical properties of the random feature models under differentially private setting provides valuable insights into differentially private learning with over-parametrized neural networks.
@@ 1.1 Contributions
The goal of this paper is to construct differentially private random features model along with theoretical guarantees. We put ourselves in the over-parametrized regime where the number of random features is greater than the number of samples. We consider the random features model obtained by solving the min-norm interpolation problem. To the best of our knowledge, it is the first work on the differentially private min-norm interpolation random features model. The construction of the DP random features model follows the output perturbation mechanism and it comes with privacy and generalization guarantees. Moreover, we study the fairness of our proposed DP random features model via the notation of excessive risk gap and of statistical parity. Our numerical experiments show that our proposed model achieves better generalization performance and does not seem to suffer unfairness. More precisely, our contributions are as follows:
@@
    We propose a differentially private over-parametrized random features model (Algorithm 1) based on a min-norm interpolator and the standard Gaussian mechanism. Our privacy guarantee relies on the sensitivity method, where we provide a novel bound on the sensitivity. Moreover, we provide a generalization error bound using the concentration results of random matrix.
@@
    We study the fairness of our proposed algorithm via the notation of excessive risk gap and of statistical parity. Following the analysis of [28] that the excessive risk gap for each group is approximated by the input norms of the elements in each group, we show that the random features method, which maps the data to a high-dimensional space where all samples have the same norms, achieves similar excessive risk gaps across all groups.
@@
    We validate our results using experiments on synthetic data and two real data sets. Our results show that our proposed method achieves better generalization performance when compared to the state-of-the-art DP random features methods in [22, 29]. We also demonstrate that the proposed DP random features model achieves fairer outputs in terms of excessive risk gap and statistical parity.
@@ 1.2 Related work
DP random features model. There is a large body of literature on privacy preserving kernel-based machine learning algorithms [19, 20, 21, 22, 29]. In [21, 20], the authors studied a private kernel Support Vector Machine (SVM) via random feature approximation of a shift-invariant kernel. In [19, 22], the authors studied the problem of privacy preserving learning using empirical risk minimization and then applied their results to produce privacy-preserving kernel-based models. Previous work assumed the empirical risk is strongly convex, and therefore a strongly convex regularization term ($\ell_{2}$ regularization for example) was added to the empirical risk function. Their theoretical guarantees depend on the choice of regularization parameter, which cannot be generalized to the interpolation regime. Our analysis, however, focuses on the interpolation regime and does not rely on the regularization term. In [29], the authors focused on the differentially private stochastic gradient descent (SGD) algorithm based on random features. The theoretical guarantees rely on technical assumptions on the kernel and a certain spectral decay rate on the second moment operator, which may be difficult to verify in practice. The obtained convergence rate depends on the sample size, which may be hard to obtain since increasing the sample size is hard in real-world applications. Our analysis does not rely on technical assumptions on the kernel and the obtained convergence rate depends on the number of parameters only.
Fairness and Privacy. Designing private fair learning algorithms has been studied in prior work [3, 2, 6, 5, 30, 28, 31]. Focusing on the classification setting, it was shown that the performance impact of privacy may be more severe on underrepresented groups, resulting in accuracy disparity [2]. In [3], the authors studied how group sample size and group clipping bias affect the impact of differential privacy in DP stochastic gradient descent and how adaptive clipping for each group helps to mitigate the disparate impact. The trade-offs between privacy and fairness were studied in [6, 28] in the regression setting. However, the trade-offs between privacy, fairness, and utility remain to be understood.
1.3 Paper Organization
The remainder of the paper is organized as follows. In Section 2, we present preliminaries on differential privacy, fairness notation, and the random features model. Our main algorithm and its theoretical guarantees are introduced in Section 3. In Section 4, we empirically explore the trade-offs achieved by our post-processing algorithm on synthetic datasets and real datasets. We conclude the paper with several future directions in Section 5.
@@ 2.1 Notation
In this paper, we let $\mathbb{R}$ be the set of all real numbers, $\mathbb{R}_{+}$ the positive real numbers, and $\mathbb{C}$ be the set of all complex numbers where $i=\sqrt{-1}$ denotes the imaginary unit. We define the set $[N]$ to be all natural numbers smaller than or equal to $N$, i.e. $\{1,2,\dots,N\}$. We use bold letters to denote vectors or matrices, and denote the identity matrix of size $n\times n$ by $\mathbf{I}_{n}$. For any two vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{d}$, the inner product is denoted by $\langle\mathbf{x},\mathbf{y}\rangle=\sum_{j=1}^{d}x_{j}\bar{y}_{j}$, where $\mathbf{x}=[x_{1},\dots,x_{d}]^{\top}$ and $\mathbf{y}=[y_{1},\dots,y_{d}]^{\top}$. For a vector $\mathbf{x}\in\mathbb{C}^{d}$, we denote by $\|\mathbf{x}\|_{p}$ the $\ell_{p}$-norm of $\mathbf{x}$. For a matrix $\mathbf{A}\in\mathbb{C}^{m\times N}$, its conjugate transpose is denoted by $\mathbf{A}^{*}$. If $\mathbf{A}\in\mathbb{C}^{m\times N}$ is full rank, then its Moore-Penrose inverse is $\mathbf{A}^{\dagger}=\mathbf{A}(\mathbf{A}\mathbf{A}^{*})^{-1}$ if $m\leq N$ or $\mathbf{A}^{\dagger}=(\mathbf{A}^{*}\mathbf{A})^{-1}\mathbf{A}^{*}$ if $m\geq N$. The (induced) $p$-norm of matrix $\mathbf{A}$ is written as $\|\mathbf{A}\|_{p}$.
@@ 2.2 Differential Privacy
2.3 Fairness Metrics
For fairness, we consider excessive risk gap and statistical parity as the fairness metrics.
Excessive Risk Gap. We first introduce excessive risk, which defines the difference between the private and non-private risk functions. Let $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be a loss function. Suppose the randomized model $f_{\theta}$ is parameterized by $\theta$, we define the risk function $L(\theta,D)$ for any given dataset $D$ as
$L(\theta,D)=\frac{1}{|D|}\sum_{(\mathbf{x}_{i},y_{i})\in D}\ell(f_{\theta}(\mathbf{x}_{i}),y_{i}).$ (1)
Then we give a formal definition of excessive risk.
Definition 3 (Excessive Risk).
Let $\theta^{\sharp}$ be the minimizer of risk function, i.e. $\theta^{\sharp}=\mathop{\mathrm{argmin}}_{\theta}L(\theta,D)$. Denote $\hat{\theta}$ the private model parameter. Then the excessive risk is defined as
where the expectation is defined over the randomness of the private mechanism.
Excessive risk is widely adopted to measure utility in private learning [32, 33]. Next we define excessive risk gap, which has been used as a metric in fairness analysis, see example [28].
Definition 4 (Excessive Risk Gap).
Denote $R(\theta,D)$ and $R(\theta,D_{a})$the population-level excessive risk and the group-level excessive risk fro sensitive group $a$, respectively. Then excessive risk gap for sensitive group $a$ is defined as
$\xi_{a}=|R(\theta)-R_{a}(\theta)|.$
Fairness is achieved when $\xi_{a}=0$ for all groups $a$. Unfairness may occur when different groups have different excessive risk gaps. Moreover, a larger excessive risk gap for a certain group implies that private estimator leads to a larger utility loss for this group.
Statistical Parity. Statistical parity requires the output distribution of the randomized regressor $f$ conditioned on each sensitive group to be similar [34]. The similarity between distributions is measured by the Kolmogorov-Smirnov distance, defined for the probability measures $p,q$ supported on $\mathbb{R}$ by
$D_{KS}(p,q)=\sup_{t\in\mathbb{R}}\left|\int_{-\infty}^{t}p(x)-q(x)dx\right|.$
Definition 5 (Statistical Parity).
A regressor $f:X\times A\to\mathbb{R}$ satisfies $\alpha$-approximate statistical parity (SP) if
$\Delta_{SP}(f):=\max_{a,a^{\prime}\in A}D_{KS}(r_{a},r_{a^{\prime}})\leq\alpha,$
where $r_{a}$ is the distribution of regressor output $f(X,A)$ conditioned on $A=a$.
2.4 Random Feature Model
Theorem 6 (Bochner [35]).
$k(\mathbf{x},\mathbf{x}^{\prime})=\int_{\mathbb{R}^{d}}\exp(i\langle\boldsymbol{\omega}_{k},\mathbf{x}-\mathbf{x}^{\prime}\rangle)d\rho(\boldsymbol{\omega})=\int_{\mathbb{R}^{d}}\exp(i\langle\boldsymbol{\omega}_{k},\mathbf{x}\rangle)\overline{\exp(i\langle\boldsymbol{\omega}_{k},\mathbf{x}^{\prime}\rangle)}d\rho(\boldsymbol{\omega}).$ (3)
The kernel function can be approximated by $k(\mathbf{x},\mathbf{x}^{\prime})\approx\langle\phi(\mathbf{x}),\phi(\mathbf{x}^{\prime})\rangle$. To produce feature map $\phi$ mapping data to finite-dimensional space $\mathbb{R}^{N}$, we can use random cosine features [36]. Setting random feature $\boldsymbol{\omega}\in\mathbb{R}^{d}$ generated from $\rho(\boldsymbol{\omega})$ and $b\in\mathbb{R}$ sampled from uniform distribution on $\pi,\pi]$, the random cosine feature is defined as $\phi(\mathbf{x},\boldsymbol{\omega},b)=\cos(\langle\boldsymbol{\omega},\mathbf{x}\rangle+b)$. For the sake of simplicity, we consider the random Fourier features in this paper. The results can be generalized to random cosine features, and therefore approximate real-valued function $f:\mathbb{R}^{d}\to\mathbb{R}$.
then problem (4) is also called the ridgeless regression problem.
@@ 3.2 Privacy Guarantee
Theorem 8 (Concentration result of Random Features Matrix).
@@
Unlike the traditional analysis on the $\ell_{2}$ sensitivity that relies on the sample size $m$ and regularization parameter, our bound depends on the number of features $N$ instead. Increasing $N$ results in smaller $\ell_{2}$-sensitivity, which leads to smaller variance of additive random noise. Hyper-parameter $\eta$ characterizes the condition number of matrix $\mathbf{A}\mathbf{A}^{*}$, which is chosen to be a fixed constant in our numerical experiments. Our theory suggests that smaller $\eta$ is favorable since it results in smaller $\ell_{2}$-sensitivity. However, this will lead to larger number of features $N$ and larger variance of random features $\sigma^{2}$, see the conditions in Theorem 8.
@@ 3.3 Generalization Error
In this section, we will derive an upper bound for the generalization error of our proposed differentially private random features model. We first define the target function class that can be approximated by random features model. Let $f:\mathbb{R}^{d}\to\mathbb{C}$ be the target function belonging to the function class
@@
where $\rho(\boldsymbol{\omega})$ is a probability distribution defined on $\mathbb{R}^{d}$. Function space $\mathcal{F}(\rho)$ is indeed a Reproducing kernel Hilbert space with associated kernel function defined in (3) via random Fourier features.
@@
$\|f-\hat{f}\|_{L^{2}(\mu)}=\left(\int_{\mathbb{R}^{d}}\left|f(\mathbf{x})-\hat{f}(\mathbf{x})\right|^{2}d\mu\right)^{1/2},$
where $\mu$ is defined by the sampling measure for the data $\mathbf{x}$. Our main result is stated in the following theorem, which characterizes the generalization error bound for differentially private random features model.
Theorem 10 (Generalization Error Bound for DP Random Features Model).
@@
and bound each term separately. The first term is the standard generalization error of random features model, which appears in recent work [38, 39]. The second term characterizes the difference between non-private and private random feature models, which can be bounded by applying the sub-exponential tail bound. ∎
n the rest of this section, we provide an error bound for each term in the right-hand side of (7). We first consider the term $\|f-f^{\sharp}\|_{L^{2}(\mu)}$, which characterizes the generalization error of the non-private random features model. Our analysis follows the structure in [38, 39], where the authors provided an upper bound for generalization error using the concentration property of random feature matrix. However, their results only hold for functions $f$ belonging to a subset of $\mathcal{F}(\rho)$. Our technical contribution is taking a different “best” random feature approximation $f^{*}$ in our analysis, which allows us to approximate all functions in $\mathcal{F}(\rho)$. Specifically, we define the best random feature approximation $f^{*}(\mathbf{x})$ as
@@
[…]
@@
Theorem 11 (Generalization Error Bound for Non-private Random Features Model.).
@@
The first term is the approximation error which measures the difference between the target function and the ”best” random features model. The second term is the estimation error. To bound the second term, we utilize the concentration property of random feature matrix (Theorem 8) as well. ∎
Now, we consider the second term $\|f^{\sharp}-\hat{f}\|_{L^{2}(\mu)}$ in (7) that characterizes the difference between the private random features model and the non-private random features model. Notice that the difference between the non-private coefficient vector $\mathbf{c}^{\sharp}$ and the private coefficient vector $\hat{\mathbf{c}}$ is indeed $\mathbf{z}\in\mathbb{R}^{N}$, which is generated from multivariate normal distribution $\mathcal{N}(0,\sigma^{2}\mathbf{I}_{N})$ with the variance $\sigma^{2}=\frac{8\ln(1.25/\delta_{p})}{N\epsilon_{p}^{2}(1-2\eta)}$ for some $\eta\in(0,1)$. In the following theorem, we present an error bound on the second term $\|f^{\sharp}-\hat{f}\|_{L^{2}(\mu)}$.
@@
$v(\mathbf{u}_{1},\dots,\mathbf{u}_{m})$ $=\|f^{\sharp}-\hat{f}\|^{2}_{L^{2}(\mu)}-\frac{1}{m}\sum_{j=1}^{m}\left|f^{\sharp}(\mathbf{u}_{j})-\hat{f}(\mathbf{u}_{j})\right|^{2}$
@@
$\leq$ $\frac{2N}{m}\|\mathbf{c}^{\sharp}-\hat{\mathbf{c}}\|_{2}^{2}\leq\frac{2N}{m}\|\mathbf{z}\|_{2}^{2}=\Delta_{v},$
@@
where the last inequality holds with probability at least $1-3\delta$ using Theorem 8. we apply Lemma 13 to estimate $\|\mathbf{z}\|_{2}^{2}$. Taking the square root for both sides leads to the desired bound. ∎
@@
Notice that $\mathbf{z}_{k}^{2}$ is a sub-exponential random variable with parameters $(\nu_{k},\alpha_{k})=(2\sigma^{2},4\sigma^{2})$, then $\|\mathbf{z}\|_{2}^{2}$ is also a sub-exponential random variable with parameters $\nu=\sqrt{\sum_{k=1}^{N}\nu_{k}^{2}}$ and $\alpha=\alpha_{k}$. Then, we apply standard sub-exponential tail bound, see [40, Proposition 2.9], to obtain the desired estimation of $\|\mathbf{z}\|_{2}^{2}$. ∎
[…]